OpenStudy (anonymous):

Find all solutions to the equation. cos2x + 2 cos x + 1 = 0

Nnesha (nnesha):

let cosx = x then deal with it as quadratic equation \[\huge\rm x^2 +2x +1 =0\]

Nnesha (nnesha):

can you u factor it ?

OpenStudy (anonymous):

So it is like (x+1)(x+)1.

OpenStudy (anonymous):

the 1 should be on the parenthesis*

Nnesha (nnesha):

yes right set it equal to zero

OpenStudy (anonymous):

It'll be -1.

Nnesha (nnesha):

x = -1 so now we can change x to cos x cos x = -1

OpenStudy (anonymous):

Is pi the only solution?

Nnesha (nnesha):

yep right or -1 has multiplicity of 2 bec cox = -1 , cosx =- 1 are solution

Nnesha (nnesha):

yes right cos x = -1 = pi

OpenStudy (anonymous):

Uhm do I have to put pi+2 pi n as the solution or just pi?

Nnesha (nnesha):

pi +2 ?

Nnesha (nnesha):

just pi

Nnesha (nnesha):

\[\pi + 2 \cancel =\pi \]

OpenStudy (anonymous):

I know that haha, but base on my textbook, "Method for Solving Trigonometric Equations Unless the equation is factorable, use substitution to get just one trigonometric function. Solve the equation for the trigonometric function. Solve for the variable in a general window of +2pin for sine and cosine and +pin for tangent. Find the specific solutions in the given interval by replacing n with integers."

OpenStudy (anonymous):

So I'm kinda confused.

OpenStudy (mertsj):

It said to find all solutions. One solution is pi. Another solution is 3pi. Another solution is 5pi... We write that like this: \[x=\pi+2\pi n\]

OpenStudy (anonymous):

@Mertsj That's what I'm talking about lol thank you for the help guys!

OpenStudy (mertsj):

yw

Nnesha (nnesha):

cool i didn't know this x = pi + 2pi n thingy